# From small eigenvalues to large cuts, and Chowla’s cosine problem

### v1

@@ 5 Spectral lower bounds for the surplus
In the first inequality, we used the fact that the scalar product of positive semidefinite matrices is nonnegative (which follows easily by considering spectral decompositions). ∎
@@ 6 Main lemmas
$W$-compression and $W$-trace. Let $W<\mathbb{R}^{n}$ be a subspace. We denote by $\Pi_{W}$ the orthogonal projection matrix onto $W$. Given an $n\times n$ symmetric matrix $M$, the $W$-compression of $M$ is the symmetric matrix
@@
From this equality, it also follows that $\operatorname{tr}_{W}(I)=\dim(W)$. We present an upper bound on the $W$-trace that will be used repeatedly in our proofs.
@@
[…]
Here, the first inequality uses that $\Pi_{W}v_{i}||\leq 1$ for every $i\in[n]$, and the second inequality is due to the Cauchy-Schwartz inequality. ∎
The importance of the $W$-compression and $W$-trace is that it allows us to analyze the contribution of the top eigenvalues of a matrix, by choosing an appropriate subspace $W$. Given a graph $G$ with adjacency matrix $A$, eigenvalues $\lambda_{1}\geq\dots\geq\lambda_{n}$ and a real number $T$, we write
@@
[…]
6.1 Main lemma - least eigenvalue version
The following lemma is the heart of our argument. It shows a curious recursive relation between the sums of the largest eigenvalues. Later, we show how to use this relation to conclude that the quadratic sum of all but the top eigenvalues are negligible. Considering the Frobenius norm, this is equivalent to saying that the adjacency matrix is well approximated by the top part of the spectral decomposition.
@@ 6.1 Main lemma - least eigenvalue version
The main idea is to analyze the identity
$A=A\circ A.$
Here, we further rewrite $A\circ A$ as
$A=(A+|\lambda_{n}|I)\circ(A+|\lambda_{n}|I)-\lambda_{n}^{2}I,$ (3)
where the advantage is that $A+|\lambda_{n}|I$ is a positive semidefinite matrix. Let $v_{1},\dots,v_{n}$ be an orthonormal basis of eigenvectors corresponding to the eigenvalues $\lambda_{1}\geq\dots\geq\lambda_{n}$. By looking at the spectral decompositions $A=\sum_{i=1}^{n}\lambda_{i}v_{i}v_{i}^{T}$ and $A+|\lambda_{n}|I=\sum_{i=1}^{n}(\lambda_{i}+|\lambda_{n}|)v_{i}v_{i}^{T}$, the identity (3) is equivalent to
$\sum_{i=1}^{n}\lambda_{i}v_{i}v_{i}^{T}=\left(\sum_{i=1}^{n}(\lambda_{i}+|\lambda_{n}|)v_{i}v_{i}^{T}\right)^{\circ 2}-\lambda_{n}^{2}I=\sum_{i,j}(\lambda_{i}+|\lambda_{n}|)(\lambda_{j}+|\lambda_{n}|)(v_{i}\circ v_{j})(v_{i}\circ v_{j})^{T}-\lambda_{n}^{2}I.$
We show that the desired inequality can be deduced by considering the $W$-traces of both sides of this identity for an appropriately chosen $W$. Define $W$ to be the subspace generated by those vectors $v_{i}\circ v_{j}$, where $\lambda_{i}$ and $\lambda_{j}$ are both at least $T$. Formally,
$W=\langle v_{i}\circ v_{j}:\lambda_{i},\lambda_{j}\geq T\rangle.$
Note that
$\dim(W)\leq N_{T}^{2}\leq\frac{S_{T}^{2}}{T^{2}}.$
Let $K=T^{2}/2n$. Then by Lemma 6.2, we have
On the other hand
[…]
We further lower bound this new sum as follows.
Claim 6.4.
[…]
Proof.
We have
[…]
Here, the right-hand-side is
$\sum_{k=1}^{n}\left(\sum_{\lambda_{i}\geq T}\lambda_{i}v_{i}(k)^{2}\right)^{2}\geq\frac{1}{n}\left(\sum_{\lambda_{i}\geq T}\sum_{k=1}^{n}\lambda_{i}v_{i}(k)^{2}\right)^{2}=\frac{S_{T}^{2}}{n},$
where the first inequality is due to the inequality between the quadratic and arithmetic mean. ∎
Using this claim, we thus proved that
Recalling that $\dim(W)\leq\frac{S_{T}^{2}}{T^{2}}$, and that the conditions of the lemma imply $\lambda_{n}^{2}\leq\frac{T^{2}}{16n}$, we can further write
$\frac{S_{T}^{2}}{n}-\lambda_{n}^{2}\dim(W)\geq\frac{S_{T}^{2}}{n}-\frac{S_{T}^{2}}{4n}=\frac{3S_{T}^{2}}{4n}.$
In conclusion, we proved that
$S_{\frac{T^{2}}{2n}}+\frac{S_{T}^{2}}{2n}\geq\operatorname{tr}_{W}(A)=\operatorname{tr}_{W}(A\circ A)\geq\frac{3S_{T}^{2}}{4n}.$
From this, the desired inequality follows. ∎
6.2 Main lemma - MaxCut version
In this section, we present a variant of the previous lemma for graphs with small MaxCut. In order to prove this lemma, we employ a similar strategy as in the proof of Lemma 6.3. However, instead of writing $A=(A+|\lambda_{n}|I)^{\circ 2}-|\lambda_{n}|I$, we write $A=(A+E)\circ(A+E)-2A\circ E-E\circ E$, where $E$ is the contribution of the negative eigenvalues. Then, most of the proof comes down to showing that if the surplus is small, then $\operatorname{tr}_{W}(A\circ E)$ and $\operatorname{tr}_{W}(E\circ E)$ are also small for an appropriately chosen $W$.
Lemma 6.5.
Let $G$ be an $n$-vertex graph such that $\operatorname{surp}^{*}(G)\leq n^{1+s}$. Let $T\geq Cn^{1-\frac{1}{24}+\frac{s}{4}}$ for some sufficiently large absolute constant $C>0$. Then
$4nS_{\frac{T^{2}}{2n}}\geq S_{T}^{2}.$
Proof.
Let $Q=\operatorname{surp}^{*}(G)$. We again analyze the identity
$A=A\circ A.$
Let $v_{1},\dots,v_{n}$ be an orthonormal basis of eigenvectors corresponding to the eigenvalues $\lambda_{1}\geq\dots\geq\lambda_{n}$. Set $E$ to be the ”negative part” of $A$, that is,
@@ 6.2 Main lemma - MaxCut version
Then we can rewrite the previous equality as
@@
Using the spectral decomposition of $A$, this is equivalent to
$\sum_{i=1}^{n}\lambda_{i}v_{i}v_{i}^{T}=\left(\sum_{\lambda_{i}>0}\lambda_{i}v_{i}v_{i}^{T}\right)\circ\left(\sum_{\lambda_{i}>0}\lambda_{i}v_{i}v_{i}^{T}\right)-2A\circ E-E\circ E,$
which can be further written as
$\sum_{i=1}^{n}\lambda_{i}v_{i}v_{i}^{T}=\sum_{\lambda_{i},\lambda_{j}>0}\lambda_{i}\lambda_{j}(v_{i}\circ v_{j})(v_{i}\circ v_{j})^{T}-2A\circ E-E\circ E.$
We bound the $W$-traces of both sides with separate methods for an appropriate subspace $W$. The terms $2A\circ E$ and $E\circ E$ constitute as error terms, for which we wish to show that their contribution to the $W$-trace is not too large.
Let $W_{0}$ be the subspace generated by those vectors $v_{i}\circ v_{j}$, where $\lambda_{i}$ and $\lambda_{j}$ are both at least $T$. Formally,
@@
The subspace $W_{0}$ is almost what we want. However, when we try to bound the contribution of the error term $\operatorname{tr}_{W_{0}}(E\circ E)$, large entries of $E$ may cause trouble. In order to overcome this, we introduce a cutoff
@@
Let $J\subset[n]$ be the set of indices $i$ such that $E_{i,i}>\beta$. Note that as $E$ is positive semidefinite, we have $\max_{i,j}|E_{i,j}|=\max_{i,i}E_{i,i}$, so $|E_{i,j}|\leq\beta$ for every $i,j\in[n]\setminus J$. We can bound the size of $J$ as follows.
@@
By Lemma 5.3, (i), we have
$\sum_{i=1}^{n}E_{i,i}=\operatorname{tr}(E)=\sum_{\lambda_{i}<0}|\lambda_{i}|\leq Q.$
From this, the claim follows immediately. ∎
@@
Consider the trace of the $W$-compressions of both sides of (5). Let $K=\frac{T^{2}}{2n}$, then using Lemma 6.2 we can upper bound the left hand side as
@@
Next, we consider the first term of the right hand side of (5), and write
[…]
@@
[…]
Using Lemma 5.4, the entries of $v_{i}$ and $v_{j}$ are bounded as $|v_{i}(k)|,|v_{j}(k)|\leq\frac{\sqrt{n}}{T}$, so we get
[…]
@@
[…]
Here, the second inequality follows by writing $\sum_{\lambda_{i},\lambda_{j}\geq T}\lambda_{i}\lambda_{j}=S_{T}^{2}$, and the last inequality follows by Claim 6.4, and writing $|J|\leq Q/\beta$.
@@
[…]
where the last equality follows by Lemma 5.3, (ii). Therefore, by Lemma 6.1, we have
[…]
Now consider $\operatorname{tr}_{W}(E\circ E)$. Let $E^{\prime}=E_{Y}$, then $E^{\prime}_{i,j}=E_{i,j}$ if $i,j\not\in J$, and $E^{\prime}_{i,j}=0$ otherwise. Therefore,
[…]
From this, using Lemma 6.1,
[…]
@@
$2\operatorname{tr}_{W}(A\circ E)+\operatorname{tr}_{W}(E\circ E)$ $\leq O(\dim(W)^{1/2}\beta n^{1/4}Q^{1/2})$
$\leq O(\dim(W)\beta n^{1/4}Q^{1/2})\leq O\left(\frac{S_{T}^{2}\beta n^{1/4}Q^{1/2}}{T^{2}}\right).$
@@
$S_{T}^{2}\left(\frac{1}{n}-O\left(\frac{Q^{3/4}n^{9/8}}{T^{3}}\right)\right)\geq\frac{3S_{T}^{2}}{4}.$
Here, the last inequality holds by our assumptions that $Q\leq n^{1+s}$ and $T\geq Cn^{1-\frac{1}{24}+\frac{s}{4}}$. In conclusion, comparing the left-hand-side and right-hand-side of (5), we get
$S_{\frac{T^{2}}{2n}}+\frac{S_{T}^{2}}{2n}\geq\operatorname{tr}_{W}(A)=\operatorname{tr}_{W}(A\circ A)\geq\frac{3S_{T}^{2}}{4n}.$
From this, the desired inequality follows. ∎

### v2

@@ 5 Spectral lower bounds for the surplus
In the first inequality, we used that $\langle v_{i}v_{i}^{T},X\rangle=v_{i}^{T}Xv_{i}\geq 0$ as $X$ is positive semidefinite. ∎
@@ 6 Main lemmas
$W$-compression and $W$-trace. Let $W<\mathbb{R}^{n}$ be a subspace. We denote by $\Pi_{W}$ the orthogonal projection matrix onto $W$. It is easy to check that $\Pi_{W}$ is symmetric. Given an $n\times n$ symmetric matrix $M$, the $W$-compression of $M$ is the symmetric matrix
@@
From this equality, it also follows that $\operatorname{tr}_{W}(I)=\dim(W)$. We present an upper bound on the $W$-trace that will be used later.
@@
$|\operatorname{tr}_{W}(M)|$ $=\left|\sum_{i=1}^{n}\mu_{i}\operatorname{tr}_{W}(v_{i}v_{i}^{T})\right|=\left|\sum_{i=1}^{n}\mu_{i}\|\Pi_{W}v_{i}\|_{2}^{2}\right|\leq\sum_{i=1}^{n}|\mu_{i}|\cdot\|\Pi_{W}v_{i}\|_{2}$
$\leq\left(\sum_{i=1}^{n}\mu_{i}^{2}\right)^{1/2}\cdot\left(\sum_{i=1}^{n}\|\Pi_{W}v_{i}\|_{2}^{2}\right)^{1/2}=\|M\|_{F}\dim(W)^{1/2}.$
Here, the first inequality uses that $\|\Pi_{W}v_{i}\|\leq 1$ for every $i\in[n]$, and the second inequality is due to the Cauchy–Schwartz inequality. ∎
The importance of the $W$-compression and $W$-trace is that it allows us to analyse the contribution of the large eigenvalues of a matrix, by choosing an appropriate subspace $W$. Given a graph $G$ with adjacency matrix $A$, eigenvalues $\lambda_{1}\geq\dots\geq\lambda_{n}$ and a positive real number $T$, we write
@@
$\operatorname{tr}_{W}(A)=\sum_{i=1}^{n}\lambda_{i}\|\Pi_{W}v_{i}\|_{2}^{2}\leq\sum_{\lambda_{i}\geq K}\lambda_{i}+K\sum_{i=1}^{n}\|\Pi_{W}v_{i}\|_{2}^{2}=S_{K}+K\dim(W).\hfill\qed$
6.1 Main lemma – smallest eigenvalue version
The following lemma is the heart of our argument. It shows a curious recursive relation between the sums of the largest eigenvalues, under the assumption that there are no very negative eigenvalues in the graph. Later, in Section˜6.3, we show how to use this relation to conclude that the quadratic sum of all but the large eigenvalues are negligible. This is, in turn, equivalent to saying that the adjacency matrix is well approximated in the Frobenius norm by the part of the spectral decomposition corresponding to the large eigenvalues, which we will discuss in more detail in Section˜8.
@@ 6.1 Main lemma – smallest eigenvalue version
$4nS_{\frac{T^{2}}{2n}}\geq S_{T}^{2}.$ (3)
The proof this lemma is prepared by a technical result, which will be used later as well.
Claim 6.4.
Let $v_{1},\dots,v_{n}$ be an orthonormal basis of eigenvectors of $A$ corresponding to the eigenvalues $\lambda_{1}\geq\dots\geq\lambda_{n}$. Then:
$\sum_{\lambda_{i},\lambda_{j}\geq T}\lambda_{i}\lambda_{j}\|v_{i}\circ v_{j}\|_{2}^{2}\geq\frac{S_{T}^{2}}{n}.$
Proof.
We can write
$\sum_{\lambda_{i},\lambda_{j}\geq T}\lambda_{i}\lambda_{j}\|v_{i}\circ v_{j}\|_{2}^{2}=\sum_{\lambda_{i},\lambda_{j}\geq T}\lambda_{i}\lambda_{j}\sum_{k=1}^{n}v_{i}(k)^{2}v_{j}(k)^{2}=\sum_{k=1}^{n}\left(\sum_{\lambda_{i}\geq T}\lambda_{i}v_{i}(k)^{2}\right)^{2}\geq\frac{1}{n}\left(\sum_{\lambda_{i}\geq T}\sum_{k=1}^{n}\lambda_{i}v_{i}(k)^{2}\right)^{2},$
where the inequality comes from the inequality between the quadratic and arithmetic mean. Since the eigenvectors are of unit length, we have $\sum_{k=1}^{n}v_{i}(k)^{2}=1$ for all $i$. Hence, the right-hand-side is equal to $\frac{1}{n}S_{T}^{2}$, finishing the proof. ∎
Proof of Lemma˜6.3.
The main idea is to analyse the identity $A=A\circ A$ using the spectral decomposition of $A$. In fact, (3) can be deduced by considering the $W$-traces of both sides of this identity for an appropriately chosen $W$. More precisely, in order to isolate the contribution of large eigenvalues, we define $W$ to be the subspace generated by those vectors $v_{i}\circ v_{j}$, where $\lambda_{i}$ and $\lambda_{j}$ are both at least $T$, i.e.
$W=\langle v_{i}\circ v_{j}:\lambda_{i},\lambda_{j}\geq T\rangle.$
Then, Lemma˜6.2 applied with $K=\frac{T^{2}}{2n}$ implies that $\operatorname{tr}_{W}(A)\leq S_{\frac{T^{2}}{2n}}+\frac{T^{2}}{2n}\dim(W)$. On the other hand, we can lower-bound $\operatorname{tr}_{W}(A\circ A)$ as follows.
Claim 6.5.
Proof.
We can rewrite $A\circ A$ as
$A\circ A$ $=(A+|\lambda_{n}|I)\circ(A+|\lambda_{n}|I)-\lambda_{n}^{2}I=\left(\sum_{i=1}^{n}(\lambda_{i}+|\lambda_{n}|)v_{i}v_{i}^{T}\right)^{\circ 2}-\lambda_{n}^{2}I$
$=\sum_{i,j}(\lambda_{i}+|\lambda_{n}|)(\lambda_{j}+|\lambda_{n}|)(v_{i}\circ v_{j})(v_{i}\circ v_{j})^{T}-\lambda_{n}^{2}I.$
Thus, the $W$-trace of $A\circ A$ can be computed as follows
$\operatorname{tr}_{W}(A\circ A)$ $=\sum_{i,j}(\lambda_{i}+|\lambda_{n}|)(\lambda_{j}+|\lambda_{n}|)\operatorname{tr}_{W}\big((v_{i}\circ v_{j})(v_{i}\circ v_{j})^{T}\big)-\lambda_{n}^{2}\operatorname{tr}_{W}(I)$
$=\sum_{i,j}(\lambda_{i}+|\lambda_{n}|)(\lambda_{j}+|\lambda_{n}|)\big\|\Pi_{W}v_{i}\circ v_{j}\big\|_{2}^{2}-\lambda_{n}^{2}\dim(W).$
$(\lambda_{i}+|\lambda_{n}|)(\lambda_{j}+|\lambda_{n}|)\big\|\Pi_{W}v_{i}\circ v_{j}\big\|_{2}^{2}\geq\lambda_{i}\lambda_{j}\|v_{i}\circ v_{j}\|_{2}^{2}.$
Also, each term in the sum is nonnegative, so if $\lambda_{i}<T$ or $\lambda_{j}<T$, we simply lower bound the contribution of $(\lambda_{i}+|\lambda_{n}|)(\lambda_{j}+|\lambda_{n}|)\|\Pi_{W}v_{i}\circ v_{j}\|_{2}^{2}$ by 0. Finally, using ˜6.4, we get
$\operatorname{tr}_{W}(A\circ A)\geq\sum_{\lambda_{i},\lambda_{j}\geq T}\lambda_{i}\lambda_{j}\|v_{i}\circ v_{j}\|_{2}^{2}-\lambda_{n}^{2}\dim(W)\geq\frac{S_{T}^{2}}{n}-\lambda_{n}^{2}\dim(W).\hfill\qed$
We now complete the proof of inequality (3). We have
$S_{\frac{T^{2}}{2n}}+\frac{T^{2}}{2n}\dim(W)\geq\operatorname{tr}_{W}(A)=\operatorname{tr}_{W}(A\circ A)\geq\frac{S_{T}^{2}}{n}-\lambda_{n}^{2}\dim(W).$
Note that $W$ has at most $N_{T}^{2}$ generators, so $\dim(W)\leq N_{T}^{2}=\frac{S_{T}^{2}}{T^{2}}.$ Finally, $|\lambda_{n}|^{2}\leq T^{2}/4n$ and thus the previous inequality implies
$S_{\frac{T^{2}}{2n}}\geq\frac{S_{T}^{2}}{n}-\Big(\frac{T^{2}}{2n}+\lambda_{n}^{2}\Big)\dim(W)\geq\frac{S_{T}^{2}}{n}-\frac{3T^{2}}{4n}\cdot\frac{S_{T}^{2}}{T^{2}}=\frac{S_{T}^{2}}{4n}.\hfill\qed$
6.2 Main lemma – MaxCut version
In this section, we present a variant of the previous lemma for graphs with small MaxCut. We employ a similar strategy as in the proof of Lemma˜6.3, but, instead of writing $A=(A+|\lambda_{n}|I)^{\circ 2}-|\lambda_{n}|I$, we write $A=(A+E)\circ(A+E)-2A\circ E-E\circ E$, where $E$ is the contribution of the negative eigenvalues. Then, most of the proof comes down to showing that if the surplus is small, then $\operatorname{tr}_{W}(A\circ E)$ and $\operatorname{tr}_{W}(E\circ E)$ are also small for an appropriately chosen $W$. For technical reasons, we prove it in terms of $\operatorname{surp}^{*}(G)$ rather than $\operatorname{surp}(G)$.
Lemma 6.6.
For every $\gamma\in(0,1/6)$, there exists a constant $C>0$ such that the following holds. Let $G$ be an $n$-vertex graph such that $\operatorname{surp}^{*}(G)\leq n^{1+\gamma}$. Then, for every $T\geq Cn^{1-\frac{1}{24}+\frac{\gamma}{4}}$,
@@ 6.2 Main lemma – MaxCut version
Let $Q=\operatorname{surp}^{*}(G)$. As in the proof of Lemma˜6.3, we analyse the identity $A=A\circ A$. Let $E$ be the “negative part” of $A$, that is,
@@
Then we can rewrite $A=A\circ A$ as
@@
The proof revolves around choosing an appropriate subspace $W$ and bounding the $W$-traces of both sides. The terms $2A\circ E$ and $E\circ E$ constitute as error terms, for which we show that their contribution to the $W$-trace is negligible.
Let $W_{0}<\mathbb{R}^{n}$ be the subspace generated by those vectors $v_{i}\circ v_{j}$, where $\lambda_{i}$ and $\lambda_{j}$ are at least $T$, i.e.
@@
The subspace $W_{0}$ is almost what we want. However, when bounding the error term $\operatorname{tr}_{W_{0}}(E\circ E)$, the large entries of $E$ have a non-negligible contribution. In order to overcome this, we introduce a cut-off
@@
Let $J\subset[n]$ be the set of indices $i$ such that $E_{i,i}>\beta$. Note that as $E$ is positive semidefinite, we have $\max_{i,j}|E_{i,j}|=\max_{i,i}E_{i,i}$, so $|E_{i,j}|\leq\beta$ for every $i,j\notin J$. Moreover, $|J|$ is small.
@@
By Lemma 5.3 (i), we have $\sum_{i=1}^{n}E_{i,i}=\operatorname{tr}(E)=\sum_{\lambda_{i}<0}|\lambda_{i}|\leq Q.$ Hence, the sum $\sum_{i=1}^{n}E_{i,i}$ contains at most $Q/\beta$ terms larger than $\beta$. ∎
@@
Consider the trace of the $W$-compressions of both sides of (4). Let $K=\frac{T^{2}}{2n}$. Then Lemma˜6.2 implies
@@
On the other hand, the term $\operatorname{tr}_{W}((A+E)\circ(A+E))$ can be lower bounded as follows.
Claim 6.8.
$\operatorname{tr}_{W}((A+E)\circ(A+E))\geq\frac{S_{T}^{2}}{n}-\frac{QS_{T}^{2}n^{2}}{\beta T^{4}}.$
Proof.
Since we have
$(A+E)\circ(A+E)=\bigg(\sum_{\lambda_{i}>0}\lambda_{i}v_{i}v_{i}^{T}\bigg)\circ\bigg(\sum_{\lambda_{i}>0}\lambda_{i}v_{i}v_{i}^{T}\bigg)=\sum_{\lambda_{i},\lambda_{j}>0}\lambda_{i}\lambda_{j}(v_{i}\circ v_{j})(v_{i}\circ v_{j})^{T},$
we can write write
$\operatorname{tr}_{W}((A+E)\circ(A+E))=\sum_{\lambda_{i},\lambda_{j}>0}\lambda_{i}\lambda_{j}\|\Pi_{W}v_{i}\circ v_{j}\|_{2}^{2}\geq\sum_{\lambda_{i},\lambda_{j}\geq T}\lambda_{i}\lambda_{j}\|\Pi_{W}v_{i}\circ v_{j}\|_{2}^{2}.$
@@
$\|\Pi_{W}(v_{i}\circ v_{j})\|_{2}^{2}=\|\Pi_{Y}(v_{i}\circ v_{j})\|_{2}^{2}=\|v_{i}\circ v_{j}\|_{2}^{2}-\sum_{k\in J}(v_{i}(k)v_{j}(k))^{2}.$
By Lemma˜5.4, the entries of $v_{i}$ and $v_{j}$ are bounded as $|v_{i}(k)|,|v_{j}(k)|\leq\frac{\sqrt{n}}{T}$, so we get
$\big\|\Pi_{W}v_{i}\circ v_{j}\big\|_{2}^{2}=\|v_{i}\circ v_{j}\|_{2}^{2}-\sum_{k\in J}(v_{i}(k)v_{j}(k))^{2}\geq\|v_{i}\circ v_{j}\|_{2}^{2}-\frac{|J|n^{2}}{T^{4}}.$
@@
$\operatorname{tr}_{W}((A+E)\circ(A+E))$ $\geq\sum_{\lambda_{i},\lambda_{j}\geq T}\lambda_{i}\lambda_{j}\left(\|v_{i}\circ v_{j}\|_{2}^{2}-\frac{|J|n^{2}}{T^{4}}\right)$
$\geq\bigg(\,\sum_{\lambda_{i},\lambda_{j}\geq T}\lambda_{i}\lambda_{j}\|v_{i}\circ v_{j}\|_{2}^{2}\,\bigg)-S_{T}^{2}\frac{|J|n^{2}}{T^{4}}\geq\frac{S_{T}^{2}}{n}-\frac{QS_{T}^{2}n^{2}}{\beta T^{4}}.$
Here, the second inequality follows by writing $\sum_{\lambda_{i},\lambda_{j}\geq T}\lambda_{i}\lambda_{j}=S_{T}^{2}$, and the last inequality follows by Claim 6.4, and writing $|J|\leq Q/\beta$. ∎
@@
$\|E\circ A\|_{F}\leq\|E\|_{F}=\bigg(\sum_{\lambda_{i}<0}\lambda_{i}^{2}\bigg)^{1/2}=O(n^{1/4}Q^{1/2}),$
where the last equality follows from Lemma˜5.3 (ii). By Lemma˜6.1, we have
$\operatorname{tr}_{W}(A\circ E)=\dim(W)^{1/2}\|E\circ A\|_{F}\leq O\big(\dim(W)^{1/2}n^{1/4}Q^{1/2}\big).$
Now consider $\operatorname{tr}_{W}(E\circ E)$. Let $E^{\prime}=E_{Y}$ be the $Y$-compression of $E$. Then $E^{\prime}_{i,j}=E_{i,j}$ if $i,j\not\in J$, and $E^{\prime}_{i,j}=0$ otherwise. Recall that $|E_{i,j}|\leq\beta$ for every $i,j\notin J$. We acquire
$\|E^{\prime}\circ E^{\prime}\|_{F}=\bigg(\,\sum_{i,j\not\in J}E_{i,j}^{4}\,\bigg)^{1/2}\leq\beta\bigg(\,\sum_{i,j\not\in J}E_{i,j}^{2}\,\bigg)^{1/2}\leq\beta\|E\|_{F}=O(\beta n^{1/4}Q^{1/2}).$
From this, using Lemma˜6.1 and that $W=\Pi_{Y}W_{0}$,
$\operatorname{tr}_{W}(E\circ E)=\operatorname{tr}_{W}(E^{\prime}\circ E^{\prime})\leq\dim(W)^{1/2}\|E^{\prime}\circ E^{\prime}\|_{F}\leq O(\dim(W)^{1/2}\beta n^{1/4}Q^{1/2}).$
@@
$2\operatorname{tr}_{W}(A\circ E)+\operatorname{tr}_{W}(E\circ E)$ $\leq O\big(\dim(W)^{1/2}\beta n^{1/4}Q^{1/2}\big)$
$\leq O\big(\dim(W)\beta n^{1/4}Q^{1/2}\big)\leq O\!\left(\frac{S_{T}^{2}\beta n^{1/4}Q^{1/2}}{T^{2}}\right).$
@@
$\operatorname{tr}_{W}(A\circ A)\geq S_{T}^{2}\left(\frac{1}{n}-O\left(\frac{Q^{3/4}n^{9/8}}{T^{3}}\right)\right)\geq\frac{3S_{T}^{2}}{4}.$
Here, the last inequality holds by our assumptions that $Q\leq n^{1+\gamma}$ and $T\geq Cn^{1-\frac{1}{24}+\frac{\gamma}{4}}$. Now, comparing the left-hand-side and right-hand-side of (4), we conclude the desired inequality by
$S_{\frac{T^{2}}{2n}}+\frac{S_{T}^{2}}{2n}$ $\geq\operatorname{tr}_{W}(A)=\operatorname{tr}_{W}(A\circ A)\geq\frac{3S_{T}^{2}}{4n}.\qed$ (5)